HUBBLE SPACE TELESCOPE DAILY REPORT #3075 PERIOD COVERED: 0000Z (UTC) 03/18/02 - 0000Z (UTC) 03/19/02 Daily Status Report as of 078/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Six Sets of ACS/WFC/HRC 8947 (Weekly Test) The Advanced Camera for Surveys (WFC and HRC) was used to perform basic tests to monitor, the read noise, the development of hot pixels and test for any source of noise in ACS CCD detectors. This program will be executed at least once a day for the entire lifetime of ACS. There were no problems reported. 1.2 Completed Two Sets of STIS/CCD 8963 (CCD Bias) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD read noise from multiple bias frames obtained in common science modes: amplifier D, gains 1 and 4 electrons/DN, and binning factors of 1x1, 1x2, 2x1, and 2x2. The read noise measurements before and after NCC turn-on are then compared, and standard calibration reference files for use in pipeline reduction of SMOV3B data are constructed. The proposal had no reported problems. 1.3 Completed Two Sets of STIS/CCD 8964 (CCD Darks) The Space Telescope Imaging Spectrograph (CCD) was used to obtain dark frames twice daily, as in a corresponding calibration program. Standard calibration reference darks are constructed and hot pixel lists are updated for use in pipeline reduction of SMOV3B data. All observations completed without incident. 1.4 Completed STIS/CCD 8906 (Hot Pixel Annealing) The Space Telescope Imaging Spectrograph (CCD) was used to measure the effectiveness of the CCD hot pixel annealing process by measuring the dark current behavior before and after annealing and by searching for any window contamination effects. In addition, CTE performance is examined by looking for traps in a low signal level flat. There were no reported problems. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 0 Successful: 0 Scheduled Re-acquisitions: 0 Successful: 0 2.2 FHST Updates: Scheduled: 0 Successful: 0 2.3 Operations Notes: Operations personnel remain in the GSFC STOCC. The NICMOS Cyro Cooler (NCC) has been recovered. Since that was accomplished, the NCC continues to operate nominally. As of 078/1130Z, the TA inlet temperature is 161.5 degK. Cooldown continues. An intercept load for the current Science Mission Specification is expected today with an expected start time of approximately 079/0000Z. COP 3.46 (Attitude Determination) is being iteratively executed in order to maintain coarse mode attitude during the period when both the FHSTs and the FGS high voltage is off for the NCS operations. At 078/0948Z, the HST coarse attitude error was 1.67 deg. Operations requests were used to command the following: NCC Circulator Start @ 077/1814z NCC Compressor Start @ 077/1817z Dump 8051 History Buffer @ 077/1833z Clear NCC flags & Turn on PCE @ 077/2028z PCE Recovery @ 077/2234z Adjust NCS CPL Reservoir SetPoint @ 077/2238z Prevent CPL Flood @ 077/2359z NCC Circulator Start @ 078/0003z NCC Compressor Start @ 078/0024z Dump 8051 History Buffer @ 078/0125z Modify 8051 History and Min/Max Buffer Collection @ 078/0146z NCS Control to PID Operate @ 078/0203z Reset Min/Max @ 078/0305z The following operations notes were executed during the day: FGS HV Monitoring during the NCS activities Coarse Attitude Determination during NCS Fill Operations FHST 2 Mounting Bracket Limit Change and Monitoring EPS Batteries 5 and 6 Current Limit Adjustment Post SM3B ROP NS-03 was utilized to perform an NSSC-1 status buffer dump at 077/2032Z. The ephemeris table was loaded at 078/0001Z per ROP DF-07A. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of Servicing Mission Orbital Verification and the gradual resumption of normal science observations and calibrations. /CAW